OpenStudy (anonymous):

Solve and check. -x - 6 = -4x - 21

OpenStudy (cristal):

Add the 4x to the left side and add the 6 to the right side to get, 3x=-15 ; divide by 3 x=-5 check: -(-5)-6=-1 -4(-5)-21= -1 -1=-1 Therefore the solution is correct.

OpenStudy (anonymous):

-x-6=-4x-21 transpose -x to the other side and 21 to the other side 21-6=-4x+x 15=-3x divide both side by -3 x=-5